# Score Operator Newton Transport

### 08 Dec 2023 v2

@@ Abstract
We propose a new approach for sampling and Bayesian computation that uses the score of the target distribution to construct a transport from a given reference distribution to the target. Our approach is an infinite-dimensional Newton method, involving an elliptic PDE, for finding a zero of a “score-residual” operator. We use classical elliptic PDE theory to prove convergence to a valid transport map. Our Newton iterates can be computed by exploiting fast solvers for elliptic PDEs, resulting in new algorithms for Bayesian inference and other sampling tasks. We identify elementary settings where score-operator Newton transport achieves fast convergence while avoiding mode collapse.
@@ 1 Introduction
  • Newton methods are efficient: we will show, empirically and in simple analytical examples (see Appendix C), that very few iterations may be required. The Newton construction also permits an existing map to be updated or fine-tuned, e.g., for perturbed scores; this is useful for applications such as Bayesian filtering.
@@
  • Elliptic differential operators instantaneously propagate information throughout the domain. Hence, our transport updates, which are elliptic PDE solutions, are intrinsically global [Evans, 2022]. As evidenced by our numerical results, our construction thus tends to avoid mode collapse, since transport updates are influenced by score values over the entire support of the distributions, including the tails.
@@ 2 Infinite-dimensional score matching
$=\Big{(}s\,(\nabla U)^{-1}-{\rm tr}\big{(}(\nabla U)^{-1}\nabla^{2}U\big{)}(\nabla U)^{-1}\Big{)}\circ U^{-1},$ (5)
@@
Note that any diffeomorphism that satisfies $\mathcal{G}(p,T)=q$ is a valid transport map in the sense that $T_{\sharp}\mu=\nu$. By integrating both sides of (5), we obtain that any $T$ that satisfies $\mathcal{G}(p,T)=q$ also satisfies, $\rho^{\nu}=k\rho^{\mu}\circ T^{-1}/|{\rm det}\nabla T|$, for an integration constant $k.$ Since the left hand side is a valid density (integrates to 1), we obtain that $k$ must be 1.
@@ 3.2 SCONE transport algorithm
In the beginning, we set the initial guess for the transport map $T_{0}$ (e.g., $T_{0}(x)={\rm Id}(x)=x$) and $p_{0}=p$, the source score. At each iteration, we solve the PDE (10) to obtain the vector field $v_{n}$. There is extensive literature on approximating solutions of elliptic PDEs using neural networks, such as physics-informed neural networks [Raissi et al., 2019], deep Ritz methods [Lu et al., 2021b, Yu et al., 2018] and Fourier neural operators [Li et al., 2020b]; as well as kernel methods and Gaussian process approximations [Owhadi and Scovel, 2019, Wendland, 2004, Schaback and Wendland, 2006, Zhang et al., 2000], and finite elements [Ciarlet, 2002, Wang and Ye, 2013]. Bespoke methods that exploit the particular structure of $\mathcal{L}(q)$ are deferred to future work (see Section 6).
@@ 3.3 Related work
The SCONE method exploits scores but does not make use of parametric spaces to define transport maps. Rather than minimizing a distance functional, we derive a generalized Newton–Raphson method on Banach spaces to construct a zero of the operator $\mathcal{R}(T)$. This approach explicitly specifies the transport map as a composition of functions that is achieved in the limit of a discrete-time dynamical system—as opposed to a continuous-time flow—on function spaces. Our SCONE transport defines, correspondingly, a discrete-time dynamical system on the space of probability densities, with the target being a fixed point.
@@ 4 Convergence proof
Here we prove the convergence of the SCONE algorithm in 1, establishing a new construction of a transport map. We do not compute explicit bounds on the error norms, $\|p_{n}-q\|$. We invoke elliptic regularity theory to establish convergence. Our Newton iterates (10) are solutions of second-order linear, elliptic systems of PDEs, as we describe below. The operator $\mathcal{L}$ is a non-self-adjoint elliptic differential operator. Let $M$ be a compact subset of $\mathbb{R}^{d}$ and $\Omega$ be an open set containing $M$. At each step, we solve $d$ second-order PDEs of the following form,
@@
Here, $\alpha=(\alpha_{1},\cdots,\alpha_{d})$ is a multi-index, $D^{\alpha}=\partial^{\alpha_{1}}_{1}\cdots\partial^{\alpha_{d}}_{d}$, such that $|\alpha|=\alpha_{1}+\cdots+\alpha_{d}\leq 2$. It turns out that our Newton update equations are strongly elliptic systems of PDEs [Morrey, 1953]. Before we recall Morrey [1953]’s notion of strong ellipticity, note that this notion guarantees similar regularity results on the solutions as those known in the more familiar single elliptic equation case (see. e.g., the textbook of Hörmander [1963] or Dyatlov [2022] for distribution results and Gilbarg and Trudinger [1977] for Schauder estimates). We say that the second-order differential operator $L(x,D)$ is strongly elliptic on $\Omega$ if the leading polynomial coefficient
$\sum_{|\alpha|=2}a^{ij}_{\alpha}(x)\xi^{\alpha}\geq C(\Omega)|\xi|^{2},\;\;$ $\forall\;\;x\in\Omega,1\leq i,j\leq d,$
$\;\xi=(\xi_{1},\cdots,\xi_{d})\neq 0,$ (14)
for some $C(\Omega)>0$. Here, $\xi^{\alpha}$ denotes the polynomial $\xi_{1}^{\alpha_{1}}\cdots\xi_{d}^{\alpha_{d}}$ and $|\xi|$ is the $\ell_{2}$ norm, i.e., $|\xi|^{2}=\sum_{i=1}^{d}|\xi_{i}|^{2}$. We repeatedly use the Schauder-type estimates from Morrey [1953] (see also Chapter 5 of Giaquinta and Martinazzi [2013]), which are estimates for the solution of second-order strongly elliptic systems in Hölder spaces.
@@
We define Hölder spaces $(\mathcal{C}^{k,\gamma}(\Omega))^{d}$ of vector-valued functions $f:=(f_{1},\ldots,f_{d})$, in such a way that $f_{i}\in\mathcal{C}^{k,\gamma}$ for each $i$. Note that, due to the (compact) embedding of Hölder spaces, $\mathcal{C}^{0,\gamma}\to\mathcal{C}^{0,\delta}$ for $\delta<\gamma$, $f\in\mathcal{C}^{k,\gamma}(\Omega))^{d}$ with $f_{i}\in\mathcal{C}^{0,\gamma_{i}}(\Omega)$ implies that $\gamma\leq\min_{i}\gamma_{i}$. In a slight abuse of notation, for a vector-valued function $f\in(\mathcal{C}^{k,\gamma}(\Omega))^{d}$, we write,
@@
Theorem 1 (Elliptic regularity for order 2 PDE systems [Morrey, 1953]).
Let $\Omega$ be a bounded and open subset of $\mathbb{R}^{d}$ with a smooth boundary. Let $L(x,D)u=f$ be a second-order strongly elliptic system of $d$ PDEs, with $L(x,D)_{ij}=\sum_{|\alpha|\leq 2}a_{\alpha}^{ij}(x)D^{\alpha}$. If the coefficients $a_{\alpha}^{ij}$ and the right hand side $f_{i}$ are in $\mathcal{C}^{s,\gamma}(\bar{\Omega})$, then, $u\in\mathcal{C}^{s+2,\gamma}(\bar{\Omega})$. In particular, for any $s\geq 0$, and $\gamma\in(0,1),$
$\|u\|_{s+2,\gamma}\leq K\|f\|_{s,\gamma},$
where $K$ only depends on $\|a_{\alpha}^{ij}\|_{s,\gamma}$ and $d$.
The above elliptic regularity and invertibility theorem is invoked repeatedly in the proof of the score-matching theorem below.
@@
The proof is an application of the inverse function theorem on Banach spaces. We set up the appropriate spaces and the operator for the application of the inverse function theorem. With $\mathcal{G}$ defined as in (5), we define a new operator $\mathcal{H}$ as
$\mathcal{H}(v)$ $:=\mathcal{G}(q,{\rm Id}+v)=\Bigg{(}q\>({\rm Id}+\nabla v)^{-1}-{\rm tr}(({\rm Id}+\nabla v)^{-1}\>\nabla^{2}v)$
@@
Let $B_{\theta}\subset\Omega$ be the $\theta$-neighborhood of $\Omega$. We define the space of $\mathcal{C}^{s}$ vector fields, $\mathcal{V}^{s,\gamma}_{\theta}:=\{v\in(\mathcal{C}^{s,\gamma}(\bar{B}_{\theta}))^{d}:v(x)\in T_{x}\mathbb{R}^{d},x\in B_{\theta}\}$ and let $B_{\theta}^{s+2,\gamma}(0)$ be a $\theta$-ball around 0 in $\mathcal{V}_{\theta}^{s+2,\gamma}$. Fix some sufficiently small $\theta>0$ such that the map $y\to y-x$ on $B_{\theta}$ is diffeomorphic on to $B_{\theta}^{s+2,\gamma}(0)$ for all $x\in\Omega$. Note that the space $\mathcal{V}_{\theta}^{s,\gamma}$ is a Banach space and the operator $\mathcal{H}:B^{s+2,\gamma}_{\theta}(0)\to(\mathcal{C}^{s,\gamma}(\bar{B_{\theta}}))^{d}$ is well-defined as a continuous operator. From definition (19), note that $\mathcal{H}(0)=q$. We can explicitly compute the first derivative of $\mathcal{H}$ at 0 to be
@@
From the above expression, it is clear that $d\mathcal{H}(0)$ is a continuous linear operator from $(\mathcal{C}^{s+2,\gamma}(\bar{B}_{\theta}))^{d}$ to $(\mathcal{C}^{s,\gamma}(\bar{B}_{\theta}))^{d}$. By checking Definition (14), we deduce that $d\mathcal{H}(0)$ is a strongly elliptic operator. Thus, applying Theorem 1, we obtain that $d\mathcal{H}(0):(\mathcal{C}^{s+2,\gamma}(\bar{B}_{\theta}))^{d}\to(\mathcal{C}^{s,\gamma}(\bar{B}_{\theta}))^{d}$ is bijective. In particular, for a fixed $q$, the $s+2$-Hölder norm of $w$ that solves $d\mathcal{H}(0)w=f$ for an $f\in\mathcal{C}^{s,\gamma}(\bar{B}_{\theta})$ is bounded above by $C\|f\|_{s,\gamma}$, for some $C$ independent of $f$. Hence, $d\mathcal{H}(0)^{-1}$ is continuous on $\mathcal{C}^{s,\gamma}(\bar{B}_{\theta})$.
@@
Thus, we can apply the inverse function theorem on Banach spaces for the operator $\mathcal{H}$. There exists an open neighborhood, $B^{s,\gamma}(q,\epsilon)$, of radius, say $\epsilon$, of $q$ in $(\mathcal{C}^{s,\gamma})^{d}$ and a continuously differentiable map $\mathcal{I}:B^{s,\gamma}(q,\epsilon)\to\mathcal{V}^{s+2,\gamma}_{\theta}$ so that $\mathcal{I}\circ\mathcal{H}(v)=v$. Thus, for any $p=\mathcal{H}(v)\in B_{s,\gamma}(q,\epsilon)$, the map $T=({\rm Id}+v)^{-1}$ is such that $\mathcal{G}(p,T)=q$. This proves $(i)$. From the continuity of $\mathcal{I}$ at $q$ in $B_{s,\gamma}(q,\epsilon)$, we can choose a $\delta_{0}>0$ such that $\|(T-{\rm Id})^{-1}\|_{s+2,\gamma}\leq\delta_{0}$. This implies that for some $\delta>0$, $\|v\|_{s+2,\gamma}\leq\delta$, hence proving (ii). ∎
@@
The above is an existence result for a transport map and is established via the inverse function theorem for the score operator. Proving the inverse function theorem via the Banach fixed point theorem both establishes the existence of the desired map $T$ and also the means to construct $T$ as a fixed point iteration of the contraction map. In the theorem below, we explicitly define such a contraction map whose fixed point is $T$. Further, the fixed point iteration of the map is equivalent to our SCONE iteration. Such an interpretation of the Newton–Raphson method as a fixed point iteration of the linearization of the given map is indeed classical in numerical analysis, when we are interested in finding zeros of a function on a finite-dimensional space. The following theorem extends this idea to infinite-dimensional spaces and serves as the convergence proof of the SCONE method.
@@
Recall the definition of $\mathcal{L}(q)$ from (8). We show in the proof of Theorem 2 that $\mathcal{L}(q)^{-1}$ is a homeomorphism between an $\theta$-neighborhood of $q$ and a $\mathcal{C}^{s+2}$ neighborhood of the identity in the space of vector fields, $B^{s,\gamma}_{\theta}(0)$. Thus, choosing $\theta>0$ appropriately, we can combine both the SCONE iteration and update steps ((10) and (11)) to define the operator,
@@
where we write $\mathcal{L}$ to indicate $\mathcal{L}(q)$ for a fixed target score $q$. It is clear that $\mathcal{J}(0)=0$ and hence 00 is a fixed point of $\mathcal{J}$. The operator $\mathcal{J}$ is smooth at 00. In particular, through direct computation, we can verify that its first derivative, $d\mathcal{J}(0)=0$ as an operator from $B^{s+2,\gamma}_{\theta}(0)$ to $\mathcal{V}^{s+2,\gamma}_{\theta}$. Applying the mean value theorem, we get,
$\|\mathcal{J}(v_{1})-\mathcal{J}(v_{2})\|\leq\|v_{1}-v_{2}\|\sup_{\eta\in(0,1)}\|d\mathcal{J}(\eta v_{1}+(1-\eta)v_{2})\|.$ (23)

### 09 Mar 2024 v3

@@ Abstract
We propose a new approach for sampling and Bayesian computation that uses the score of the target distribution to construct a transport from a given reference distribution to the target. Our approach is an infinite-dimensional Newton method, involving an elliptic PDE, for finding a zero of a “score-residual” operator. We prove sufficient conditions for convergence to a valid transport map. Our Newton iterates can be computed by exploiting fast solvers for elliptic PDEs, resulting in new algorithms for Bayesian inference and other sampling tasks. We identify elementary settings where score operator Newton transport achieves fast convergence while avoiding mode collapse.
@@ 1 Introduction
  • Newton methods are efficient: we will show, empirically and in simple analytical examples (see Appendix B), that very few iterations may be required. The Newton construction also permits an existing map to be updated or fine-tuned, e.g., for perturbed scores; this is useful for applications such as Bayesian filtering.
@@
  • Elliptic differential operators instantaneously propagate information throughout the domain. Hence, our transport updates, which use elliptic PDE solutions, are intrinsically global [Evans, 2022]. As evidenced by our numerical results, our construction thus tends to avoid mode collapse, since transport updates are influenced by score values over the entire support of the distributions, including the tails.
@@ 2 Infinite-dimensional score matching
$=\Big{(}s\,(\nabla U)^{-1}-{\rm tr}\big{(}(\nabla U)^{-1}$ (5)
$\nabla^{2}U\big{)}(\nabla U)^{-1}\Big{)}\circ U^{-1},$
@@
Note that any diffeomorphism that satisfies $\mathcal{G}(p,T)=q$ is a valid transport map in the sense that $T_{\sharp}\mu=\nu$. By integrating both sides of (5), we obtain that any $T$ that satisfies $\mathcal{G}(p,T)=q$ also satisfies, $\rho^{\nu}=k(\rho^{\mu}/|{\rm det}\nabla T|)\circ T^{-1}$, for an integration constant $k.$ Since the left hand side is a valid density (integrates to 1), we obtain that $k$ must be 1.
@@ 3.2 SCONE transport algorithm
In the beginning, we set the initial guess for the transport map $T_{0}$ (e.g., $T_{0}(x)={\rm Id}(x)=x$) and $p_{0}=p$, the source score. At each iteration, we solve the PDE (10) to obtain the vector field $v_{n}$. There is extensive literature on approximating solutions of PDEs using neural networks, such as physics-informed neural networks [Raissi et al., 2019], deep Ritz methods [Lu et al., 2021b, Yu et al., 2018] and Fourier neural operators [Li et al., 2020b]; as well as kernel methods and Gaussian process approximations [Owhadi and Scovel, 2019, Wendland, 2004, Schaback and Wendland, 2006, Zhang et al., 2000], and finite elements [Ciarlet, 2002, Wang and Ye, 2013]. Bespoke methods that exploit the particular structure of $\mathcal{L}(q)$ are deferred to future work (see Section 6).
@@ 3.3 Related work
The SCONE method exploits scores but does not make use of parametric spaces to define transport maps. Rather than minimizing a distance functional, we derive a generalized Newton–Raphson method on Banach spaces to construct a zero of the operator $\mathcal{R}(T)$. This approach specifies the transport map as a composition of functions that is achieved in the limit of a discrete-time dynamical system—as opposed to a continuous-time flow—on function spaces. Our SCONE transport defines, correspondingly, a discrete-time dynamical system on the space of probability densities, with the target being a fixed point.
@@ 4 Convergence proof
Here we prove the convergence of the SCONE algorithm in 1, establishing a new construction of a transport map. We do not compute explicit bounds on the error norms, $\|p_{n}-q\|$. We invoke elliptic regularity theory to establish convergence. Our Newton iterates (10) yield second-order linear, elliptic PDEs, as we describe below. Let $M$ be a compact subset of $\mathbb{R}^{d}$ and $\Omega$ be an open set containing $M$. At each step, we solve $d$ second-order PDEs of the following form,
@@
Here, $\alpha=(\alpha_{1},\cdots,\alpha_{d})$ is a multi-index, $D^{\alpha}=\partial^{\alpha_{1}}_{1}\cdots\partial^{\alpha_{d}}_{d}$, such that $|\alpha|=\alpha_{1}+\cdots+\alpha_{d}\leq 2$. Suppose we parameterize the solutions $v_{n}$ of the system as the gradient $\nabla\phi_{n}$ of some differentiable function $\phi_{n}:\Omega\to\mathbb{R}.$ Then, substituting into (10), we obtain the following equation for the scalar function $\phi_{n},$
$\nabla(\nabla\cdot\nabla\phi_{n})+\nabla(\nabla\phi_{n}\cdot q)=p_{n}-q.$
Integrating this equation, we find that our Newton iterates, when $v_{n}=\nabla\phi_{n}$ satisfies,
$\nabla\cdot\nabla\phi_{n}+\nabla\phi_{n}\cdot q=\log(\rho^{\mu}_{n}/\rho^{\nu})+C,$ (14)
where $C$ is an integration constant and $\rho^{\mu}_{n}$ is the density of the pushforward measure, $T_{n\sharp}\mu.$ The operator, $\mathcal{P}=\nabla.\nabla+q\cdot\nabla,$ is an elliptic operator, which ensures the well-posedness of solutions to (14). We recall the ellipticity condition of an operator, which says that the highest-order (principal) symbol is coercive (see e.g., the textbook of Hörmander [1963] or Dyatlov [2022] for notes and Gilbarg and Trudinger [1977] for Schauder estimates).
@@
We define Hölder continuous cotangent vector fields of the form $v=\nabla\phi\in D^{*}\bar{\Omega}$ for some differentiable function $\phi,$ where $D^{*}\bar{\Omega}$ is used to denote the cotangent bundle (rather than the usual $T^{*}\bar{\Omega},$ to avoid confusion with transport maps $T$). If $\phi\in\mathcal{C}^{k,\gamma}(\Omega)$, then, we have that the components of $v$ (interpreted as scalar functions) are $\mathcal{C}^{k-1,\gamma}(\Omega).$ Note that, due to the (compact) embedding of Hölder spaces, $\mathcal{C}^{0,\gamma}\to\mathcal{C}^{0,\delta}$ for $\delta<\gamma$, $f\in(\mathcal{C}^{k,\gamma}(\Omega))^{d}$ with components $f_{i}\in\mathcal{C}^{0,\gamma_{i}}(\Omega)$ implies that $\gamma\leq\min_{i}\gamma_{i}$. In a slight abuse of notation, for a vector-valued function $f\in(\mathcal{C}^{k,\gamma}(\Omega))^{d},$ including cotangent vector fields, we write,
@@
We use Schauder estimates of the type below [Gilbarg and Trudinger, 1977] from classical elliptic PDE theory.
Theorem 1.
Let $\Omega$ be a bounded and open subset of $\mathbb{R}^{d}$ with a smooth boundary. Let $L(x,D)u=f$ be a second-order strongly elliptic system, with $L(x,D)=\sum_{|\alpha|\leq 2}a_{\alpha}(x)D^{\alpha}$, and zero Dirichlet boundary conditions. If the coefficients $a_{\alpha}$ and the right hand side $f$ are in $\mathcal{C}^{s,\gamma}(\bar{\Omega})$, then, $u\in\mathcal{C}^{s+2,\gamma}(\bar{\Omega})$. In particular, for any $s\geq 0$, and $\gamma\in(0,1),$
$\|u\|_{s+2,\gamma}\leq K(\|f\|_{s,\gamma}+\|u\|_{s}),$
where $K$ only depends on $\|a_{\alpha}\|_{s,\gamma}$ and $d$.
@@
Define
$\mathcal{H}(v)$ $:=\mathcal{G}(q,{\rm Id}+v)=\Bigg{(}q\>({\rm Id}+\nabla v)^{-1}-$
${\rm tr}(({\rm Id}+\nabla v)^{-1}\>\nabla^{2}v)$
@@
From definition (19), note that $\mathcal{H}(0)=q$. We can explicitly compute the first derivative of $\mathcal{H}$ at 0 to be
@@
We can also deduce that $d\mathcal{H}(0)\nabla\phi$ is equivalent to $-\nabla\mathcal{P}\phi,$ where $\mathcal{P}=\nabla\cdot\nabla+q\cdot\nabla$ is an elliptic operator, which is Fredholm on $\mathcal{C}^{s+2,\gamma}(\bar{\Omega})$. Let $C^{s,\gamma}(\bar{\Omega})$ denote the quotient space of $\mathcal{C}^{s,\gamma}(\bar{\Omega})$ corresponding to the equivalence relation $f\sim g$ if $\nabla f(x)=\nabla g(x),$ at all $x\in\bar{\Omega}.$ Correspondingly, we define a closed subspace of $\mathcal{C}^{s}$ cotangent vector fields, $\mathcal{V}^{s,\gamma}:=\{x\to v(x)=\nabla\phi(x)\in D_{x}^{*}\bar{\Omega},\phi\in{C}^{s+1,\gamma}(\bar{\Omega}),x\in\Omega\},$ with norm $\|\nabla\phi\|_{*}=\|\nabla\phi\|_{s,\gamma}$ and let $B_{\theta}^{s+2,\gamma}(0)$ be a $\theta$-ball around 0 in $\mathcal{V}^{s+2,\gamma}$. The element, $\phi,$ in a sufficiently small set around the constant element in $C^{s+3,\gamma}(\bar{\Omega})$ can identify an element, $\nabla\phi,$ of $B_{\theta}^{s+2,\gamma}(0).$ We note that the operator $\mathcal{H}:B^{s+2,\gamma}_{\theta}(0)\to(\mathcal{C}^{s,\gamma}(\bar{\Omega}))^{d}$ is well-defined as a continuous operator.
When $d\mathcal{H}(0)$ is defined on $\mathcal{V}^{s+2,\gamma},$ and using Theorem 1, we know that its kernel only contains the zero element of $\mathcal{V}^{s+2,\gamma},$ which corresponds to $\mathcal{P}\phi=$ const $\in C^{s,\gamma}(\bar{\Omega})$. Thus, we obtain that $d\mathcal{H}(0):\mathcal{V}^{s+2,\gamma}\to\mathcal{V}^{s,\gamma}$ is bijective. In particular, for a fixed $q$, the $s+2$-Hölder norm of $w$ that solves $d\mathcal{H}(0)w=f$ for an $f\in\mathcal{V}^{s,\gamma}$ is bounded above, by Theorem 1. Hence, $d\mathcal{H}(0)^{-1}$ is continuous on $\mathcal{V}^{s,\gamma}$.
@@
Thus, we can apply the inverse function theorem for $\mathcal{H}$. There exists an open neighborhood, $B^{s,\gamma}(q,\epsilon)$, of radius, say $\epsilon$, of $q$ in $\mathcal{V}^{s,\gamma}$ and a continuously differentiable map $\mathcal{I}:B^{s,\gamma}(q,\epsilon)\to\mathcal{V}^{s+2,\gamma}$ so that $\mathcal{I}\circ\mathcal{H}(v)=v$. Thus, for any $p=\mathcal{H}(v)\in B_{s,\gamma}(q,\epsilon)$, the map $T=({\rm Id}+v)^{-1}$ is such that $\mathcal{G}(p,T)=q$. This proves $(i)$. From the continuity of $\mathcal{I}$ at $q$ in $B_{s,\gamma}(q,\epsilon)$, we can choose a $\delta_{0}>0$ such that $\|(T-{\rm Id})^{-1}\|_{s+2,\gamma}\leq\delta_{0}$. This implies that for some $\delta>0$, $\|v\|_{s+2,\gamma}\leq\delta$, hence proving (ii). ∎
The above is an existence result for a transport map and is established via the inverse function theorem for the score operator. It is important to note that even though the result is local (that is, for nearby probability densities), uniqueness cannot still be established for the transport map. This is because, in the above proof, the operator $d\mathcal{H}(0)$ has a non-empty kernel on function spaces containing functions of the form, $\nabla\phi,$ for some $\phi\in\mathcal{C}^{s+2,\gamma}(\bar{\Omega}).$ For any function $\phi$ such that $\mathcal{P}\phi=f+{\rm const},$ $v=\nabla\phi$ solves $d\mathcal{H}(0)v=\nabla f,$ and therefore $d\mathcal{H}(0)$ is not injective. In other words, we have not shown that there is only one transport map between a given source and target density, even when they are close to each other. We have defined quotient spaces on which to define $d\mathcal{H}(0)$ to make it invertible, using isomorphism theorems for vector spaces.
@@
Proving the inverse function theorem via the Banach fixed point theorem both establishes the existence of the desired map $T$ and also the means to construct $T$ as a fixed point iteration of the contraction map. In the theorem below, we explicitly define such a contraction map whose fixed point is $T$. Further, the fixed point iteration of the map is equivalent to our SCONE iteration. Such an interpretation of the Newton–Raphson method as a fixed point iteration of the linearization of the given map is indeed classical in numerical analysis, when we are interested in finding zeros of a function on a finite-dimensional space. The following theorem extends this idea to infinite-dimensional spaces and serves as the convergence proof of the SCONE method.
@@
Recall the definition of $\mathcal{L}(q)$ from (8). Note that $\mathcal{L}(q)$ is not invertible on $(\mathcal{C}^{s,\gamma}(\bar{\Omega}))^{d},$ and $v_{n}$ in the statement of the theorem refers to any cotangent vector field $v_{n}=\nabla\phi_{n}$ such that $\mathcal{L}(q)v_{n}=p_{n}-q.$ We show in the proof of Theorem 2 that $\mathcal{L}(q)^{-1}$ is a homeomorphism between a $\theta$-neighborhood of $q$ in $\mathcal{V}^{s,\gamma}$ (see the proof of 2 for the definition of this space) and a $\mathcal{V}^{s+2,\gamma}$ neighborhood of zero. Thus, choosing $\theta>0$ sufficiently small, we can combine both the SCONE iteration and update steps ((10) and (11)) to define the operator,
@@
where we write $\mathcal{L}$ to indicate $\mathcal{L}(q)$ for a fixed target score $q$. It is clear that $\mathcal{J}(0)=0$ and hence 00 is a fixed point of $\mathcal{J}$. The operator $\mathcal{J}$ is smooth at 00. In particular, through direct computation, we can verify that its first derivative, $d\mathcal{J}(0)=0$ as an operator from $B^{s+2,\gamma}_{\theta}(0)$ to $\mathcal{V}^{s+2,\gamma}$. Applying the mean value theorem, we get,
$\|\mathcal{J}(v_{1})-\mathcal{J}(v_{2})\|$ $\leq\|v_{1}-v_{2}\|$
$\sup_{\eta\in(0,1)}\|d\mathcal{J}(\eta v_{1}+(1-\eta)v_{2})\|.$ (23)